# Agentic AI for 6G: A New Paradigm for Autonomous RAN Security Compliance

### v1

@@ I Introduction
The telecommunications industry is rapidly adopting artificial intelligence (AI)–driven solutions to manage the growing complexity and scale of next-generation mobile networks. Traditional machine learning (ML) and deep learning (DL) techniques have demonstrated strong performance in tasks such as resource allocation, network optimization, and anomaly detection. However, next-generation networks are increasingly evolving toward an intent-driven paradigm, wherein operators specify high-level goals rather than explicit configuration parameters. In such environments, existing ML and DL methods often lack the contextual understanding needed to interpret operator intents. To address these emerging challenges, recent research has turned to large language models (LLMs) as a promising new paradigm, offering advanced reasoning, interpretability, and contextual awareness that can support intelligent, adaptive network management and automation. Yet, despite their versatility, LLMs trained on generic datasets often lack the domain-specific knowledge and technical precision required for telecommunications applications. To fully harness their potential, there is a growing need for telecommunications foundation models trained on industry-relevant datasets and aligned with standardization requirements [maatouk2024large, zhou2024large].
Among the emerging applications of LLMs in telecommunications is network security. Researchers are increasingly investigating how LLMs can enhance threat modeling, vulnerability assessment, and anomaly detection. In the context of 5G and 6G networks, efforts have integrated LLMs into the STRIDE threat modeling framework to automate threat classification and recommend appropriate mitigations [abdulghaffar2025llms]. In parallel, LLMs have been proposed as autonomous reasoning engines capable of analyzing security risks, correlating multi-layer telemetry, and dynamically adapting security policies in real time [nguyen2024large]. Recent studies have further applied LLM-based approaches within Open RAN (O-RAN) and AI-RAN environments, demonstrating their potential for explainable intrusion detection [chatzimiltis2025xaillm]. Despite the advanced reasoning capabilities inherent in LLMs, their standalone use faces fundamental limitations that restrict their effectiveness for deployment in real-world telecommunication environments. A primary constraint is the static nature of their pre-trained knowledge, which limits their ability to incorporate real-time network states, evolving standards, and domain-specific data. Furthermore, LLMs are fundamentally reactive, requiring explicit human prompting for every action, which limits their suitability for dynamic environments demanding autonomous, time-critical decisions.
@@
[dev2025adv] 6G architectures and V2X optimization via multi-agent reasoning
[sharma2025mobillm] O-RAN threat mitigation using multi-agent LLMs with FiGHT-guided closed-loop responses
[nezami2025standardization] Standardization of GenAI-driven agentic architectures for RAN
[elkael2025agentran] Intent-driven autonomous RAN control via multi-layer LLM-based agents
[salama2025edge] Edge-based autonomous RAN optimisation using persona-driven agents with predictive LSTM tools
[wu2025llm] Adaptive O-RAN slice resource optimization using an LLM-driven agentic xApp
[chatzistefanidis2025mxaiagenticobservabilitycontrol] Intent-driven O-RAN observability and configuration control via multi-agent planning and telemetry retrieval
@@
Consequently, recent works have begun to investigate agentic AI across the 6G landscape, from architectural design to security and resource optimization. At the architectural level, recent research positions agentic AI as a key building block for AI-native 6G networks, enabling intent-driven orchestration, multi-domain automation, and streamlined control/user plane interactions [dev2025adv]. The same paradigm is gaining traction in the security domain, where agentic frameworks are used for closed-loop threat mitigation within O-RAN environments [sharma2025mobillm]. Standardization efforts are converging toward this direction, with emerging work defining how agentic intelligence should be embedded into O-RAN and AI-RAN architectures [nezami2025standardization]. In parallel, architectural proposals such as AgentRAN illustrate how hierarchical teams of LLM-based agents can operationalize intent-driven, multi-timescale control across open 6G networks [elkael2025agentran]. Complementary system-level studies further demonstrate that agentic AI can support resource optimization [salama2025edge], enhance RAN resilience [wu2025llm], and enable distributed observability and coordinated control across heterogeneous RAN components [chatzistefanidis2025mxaiagenticobservabilitycontrol].
Despite this rapid progress, current efforts remain largely focused on optimization and threat mitigation, leaving broader aspects of security assurance comparatively unexplored. In line with ongoing standardization efforts across ITU, 3GPP, and the O-RAN Alliance [itu2025genaitelecom, oran2025genai], our work positions agentic AI as a key enabler for trustworthy, explainable, and autonomous security assurance, an essential capability for customer-centric, AI-native 6G networks. As RAN architectures evolve, ensuring that network components and services conform to changing security requirements necessitates both specification-aware intelligence (i.e., documentation-based compliance) and runtime observability (i.e., dynamic, evidence-driven monitoring). Traditional compliance methods, which rely mainly on static documentation and offline analysis, are insufficient in environments where configurations evolve continuously through intelligent applications. Meeting this challenge calls for a unified compliance model that spans both compliance by design and compliance by evidence [salman2024compliance]. Such a model requires systems not only to demonstrate conformance to formal standards specifications but also to provide contextual, verifiable justifications for their behaviour at runtime.

### v2

@@ I Introduction
The telecommunications industry is rapidly adopting artificial intelligence (AI)–driven solutions to manage the growing complexity and scale of next-generation mobile networks. Traditional machine learning (ML) and deep learning (DL) techniques have demonstrated strong performance in tasks such as resource allocation, network optimization, and anomaly detection. However, next-generation networks are increasingly evolving toward an intent-driven paradigm, wherein operators specify high-level goals rather than explicit configuration parameters. In such environments, existing ML and DL methods often lack the contextual understanding needed to interpret operator intents. To address these emerging challenges, recent research has turned to large language models (LLMs) as a promising new paradigm, offering advanced reasoning, interpretability, and contextual awareness that can support intelligent, adaptive network management and automation. Yet, despite their versatility, LLMs trained on generic datasets often lack the domain-specific knowledge and technical precision required for telecommunications applications. To fully harness their potential, there is a growing need for telecommunications foundation models trained on industry-relevant datasets and aligned with standardization requirements [6, 15].
Among the emerging applications of LLMs in telecommunications is network security. Researchers are increasingly investigating how LLMs can enhance threat modeling, vulnerability assessment, and anomaly detection. In the context of 5G and 6G networks, efforts have integrated LLMs into the STRIDE threat modeling framework to automate threat classification and recommend appropriate mitigations [1]. In parallel, LLMs have been proposed as autonomous reasoning engines capable of analyzing security risks, correlating multi-layer telemetry, and dynamically adapting security policies in real time [8]. Recent studies have further applied LLM-based approaches within Open RAN (O-RAN) and AI-RAN environments, demonstrating their potential for explainable intrusion detection [2]. Despite the advanced reasoning capabilities inherent in LLMs, their standalone use faces fundamental limitations that restrict their effectiveness for deployment in real-world telecommunication environments. A primary constraint is the static nature of their pre-trained knowledge, which limits their ability to incorporate real-time network states, evolving standards, and domain-specific data. Furthermore, LLMs are fundamentally reactive, requiring explicit human prompting for every action, which limits their suitability for dynamic environments demanding autonomous, time-critical decisions.
@@
[4] 6G architectures and V2X optimization via multi-agent reasoning
[12] O-RAN threat mitigation using multi-agent LLMs with FiGHT-guided closed-loop responses
[7] Standardization of GenAI-driven agentic architectures for RAN
[5] Intent-driven autonomous RAN control via multi-layer LLM-based agents
[10] Edge-based autonomous RAN optimisation using persona-driven agents with predictive LSTM tools
[14] Adaptive O-RAN slice resource optimization using an LLM-driven agentic xApp
[3] Intent-driven O-RAN observability and configuration control via multi-agent planning and telemetry retrieval
@@
Consequently, recent works have begun to investigate agentic AI across the 6G landscape, from architectural design to security and resource optimization. At the architectural level, recent research positions agentic AI as a key building block for AI-native 6G networks, enabling intent-driven orchestration, multi-domain automation, and streamlined control/user plane interactions [4]. The same paradigm is gaining traction in the security domain, where agentic frameworks are used for closed-loop threat mitigation within O-RAN environments [12]. Standardization efforts are converging toward this direction, with emerging work defining how agentic intelligence should be embedded into O-RAN and AI-RAN architectures [7]. In parallel, architectural proposals such as AgentRAN illustrate how hierarchical teams of LLM-based agents can operationalize intent-driven, multi-timescale control across open 6G networks [5]. Complementary system-level studies further demonstrate that agentic AI can support resource optimization [10], enhance RAN resilience [14], and enable distributed observability and coordinated control across heterogeneous RAN components [3].
Despite this rapid progress, current efforts remain largely focused on optimization and threat mitigation, leaving broader aspects of security assurance comparatively unexplored. In line with ongoing standardization efforts across ITU, 3GPP, and the O-RAN Alliance [13, 9], our work positions agentic AI as a key enabler for trustworthy, explainable, and autonomous security assurance, an essential capability for customer-centric, AI-native 6G networks. As RAN architectures evolve, ensuring that network components and services conform to changing security requirements necessitates both specification-aware intelligence (i.e., documentation-based compliance) and runtime observability (i.e., dynamic, evidence-driven monitoring). Traditional compliance methods, which rely mainly on static documentation and offline analysis, are insufficient in environments where configurations evolve continuously through intelligent applications. Meeting this challenge calls for a unified compliance model that spans both compliance by design and compliance by evidence [11]. Such a model requires systems not only to demonstrate conformance to formal standards specifications but also to provide contextual, verifiable justifications for their behaviour at runtime.